OpenStudy (anonymous):

(sinx-cosx)(sinx+cosx)/sin^2x-1/sec^2x

OpenStudy (abb0t):

And what the heck are we supposed to do with this nonsense?

OpenStudy (anonymous):

Simplify.

OpenStudy (aum):

\( \Large \frac{(\sin x - \cos x)(\sin x + \cos x)}{\left ( \sin^2 x - \frac{1}{\sec^2 x}\right)} = \Large \frac{(\sin x - \cos x)(\sin x + \cos x)}{\left ( \sin^2 x - \cos^2 x\right)} = \\ \Large \Large \frac{(\sin x - \cos x)(\sin x + \cos x)}{(\sin x + \cos x)(\sin x - \cos x)} = 1 \)

OpenStudy (aum):

Since \(a^2 - b^2 = (a+b)(a-b)\)

OpenStudy (anonymous):

Ok